x-3y=13
5x+3y=2
elimination

Answer by njusa(12)

x-3y=13....1
5x+3y=2...2
1+2:
6x=15
x=5/2
substitue x into the solution:
5/2 -3y = 13
-3y = 21/2
y= -7/2
Mukhi